# The taggedPBC: Annotating a massive parallel corpus for crosslinguistic investigations

### v2

@@ 3 Results
The final taggedPBC dataset contains tagged data for 1,942 modern languages (1,944 if you include the 2 conlangs with NT translations: Esperanto and Klingon). This represents roughly 25% of the world’s languages from 155 distinct families and 78 single languages or isolates. The number of languages with more than 1,800 sentences/verses in the dataset is 1,853 (see Table 1), and the fewest number of verses is represented by Nafusi (ISO 639-3: jbn) with 729 verses. The automatic tagging method outlined above results in a statistically robust number of unique arguments (combining nouns, proper nouns, pronouns) and predicates (combining verbs and auxiliaries) for all 1,942 languages (mean: 699 arguments, 466 predicates), with the fewest number of unique arguments being tagged for Nugunu (ISO 639-3: yas; 76) and the fewest number of unique predicates being tagged for Rapanui (ISO 639-3: rap; 57). If we compare the number of unique nouns and verbs, the fewest unique nouns are identified for ut-Hun (ISO 639-3: uth; 59) and the fewest unique verbs for Bambalang (ISO 639-3: bmo; 54).
@@ 3.1 Validating tag accuracy
There are two means by which we can validate the automated tags provided by the IBM2 word alignment method. The first involves assessing against other automatic taggers developed for high resource languages. Here we want to see that the method used for tagging these languages aligns with existing language-specific taggers at a level above chance. The second means of assessment involves comparing the words and their tags with hand-tagged data in gold standard corpora. Here, if we find exact matches (in both word form and tag) between the taggedPBC corpora and those in hand-tagged corpora, we can be confident that other words in the taggedPBC with the same part of speech do in fact belong to that word class. For the following evaluations I focus strictly on the word classes of nouns and verbs, as these are most well-represented in the data.
3.1.1 Validation of tags via automated POS-taggers
POS-taggers have been developed for high-resource languages, and the current state of the art is represented by the SpaCy [^26] https://spacy.io/ and Trankit [^27] http://nlp.uoregon.edu/trankit text processing libraries. These libraries have supervised POS-taggers for 16 and 31 languages respectively (excluding languages with non-roman scripts). For this comparison, each of the trained taggers in these libraries was used to tag the reduced PBC verses for the respective language and then the specific tags for arguments (NOUN, PROPN, PRON) and predicates (VERB, AUX) were compared with the same verses tagged via the word alignment method. [^28] I should note here, first, that both libraries use a variation of the universal tag set adopted by the UDT project and my methodology, as outlined in footnote 25. While there are some differences, they apply to less common word classes and do not affect arguments and predicates, which were the focus of this comparison. Second, these taggers were not developed using Biblical texts, which reduces the overall likelihood of correspondence, and makes the actual result more striking. Specifically, each word-alignment tagged verse was stripped of tags and re-tagged by the respective libraries. Then the two sets of tagged verses (between the library and the word-alignment method) were compared and the number of matching arguments/predicates was divided by the total number of arguments/predicates identified by the respective library in a given sentence. This was then averaged across all verses to provide an agreement percentage between the translation method and the respective library for the argument and predicate tags. [^29] See code at the linked repository for a more detailed breakdown of correspondence between individual POS tags in the argument and predicate classes. Correspondence with argument and predicate tags between the word alignment method and these pos-taggers was on average above 75% for both SpaCy (Table 2; arguments: 0.84; predicates: 0.76) and Trankit (Table 3; arguments: 0.84; predicates: 0.77).
@@ 3.1.1 Validation of tags via automated POS-taggers
This comparison informs us that the translation method agrees with existing SOTA POS-taggers more than 75% of the time, which suggests that this methodology is relatively robust at transferring tags from English to other languages. At least for tagging major word classes, it aligns pretty well with automatic taggers individually developed for specific languages.
3.1.2 Validation of tags via gold standard corpora
The most comparable dataset with tagged corpus data, as noted above, is the UDT. As of the v2.14 release, there were 160 languages represented in this dataset. Of this data, 106 languages contain more than 200 hand-tagged sentences, and 86 of these languages are also present in the taggedPBC. Given that the UDT contains a mixture of different genres of texts for each language, this provides a good validation of the IBM2 pos-tagging methodology outlined above. If the method used to tag the PBC results in word/tag pairs that exactly match with both words and tags in the UDT for a given language (i.e. for both datasets, in Yakut [ISO 639-3: sah] kiliebi is tagged as ‘NOUN’ and kel is tagged as ‘VERB’), we can surmise that other words in the PBC with the same tags are highly likely to belong to the same class.
Despite the PBC being a rather different genre of texts than those that make up the UDT corpora, this is indeed the case: there are words with the exact same form and tag for every one of the 86 languages shared by the two corpora where the UDT has at least 200 sentences. On average, each pair of corpora shares 137 nouns and 89 verbs as exact matches. In fact, there is only one language with one or fewer verbs that match in form and POS tag: Urdu (ISO 639-3: urd) which shares one verbal form (kin) between the two tagged corpora. When we investigate this more closely, it turns out that in the Urdu UDT corpus there are no diacritics to mark vowels, whereas the Urdu PBC corpus contains diacritics that allow for insertion of vowels in the romanized script. The exact matches and code to identify these correspondences can be found at the linked repository.
The two points of comparison (between automated taggers on the one hand and hand-tagged corpora on the other) indicate that the method used for tagging parts of speech for low-resource languages in the PBC is a valid approach, given that it largely aligns with POS taggers for high-resource languages and at the same time identifies actual nouns and verbs as found in hand-tagged data. However, “the proof is in the pudding”: does the data in the taggedPBC lend itself to actual correlations with observed properties of language? To assess this, we now turn to observations of word order that can be extracted from this dataset.
@@ 3.2 Testing features extracted from the dataset: word order
By “word order” most researchers refer to the relative order of referential arguments (nouns, proper nouns, pronouns) versus predicates (verbs, auxiliaries), such that it is possible to talk about “verb-initial” vs “verb-final” or “verb-medial” languages, as well as grammatical constructs like Subject and Object in relation to the position of the Verb. In the current investigation I focus on intransitive word order, which is concerned with a predicate that requires a single core argument, resulting in three possible word orders: “verb final” (SV), “verb initial” (VS), and “free” (no dominant order). In order to identify these relations in corpora I use the following procedure for each language in the taggedPBC corpus:
@@
    Make lists of all the verses in the corpus, filtering out verses/sentences which do not contain both an argument and a predicate.
@@
    In the resulting set, verses/sentences with an argument occurring first are considered to be “noun first”, while those with a predicate occurring first are considered “verb first”.
@@
    This gives a count of noun-first and verb-first verses/sentences for each language.
@@
    Turn this into the “N1 ratio” by dividing the total number of noun-first verses/sentences by the total number of verb-first verses/sentences.
@@
To test the validity of this measure I compared the derived values with known expert judgements of intransitive word order (SV, VS, or “free”) in three typological databases (Figure 1): the World Atlas of Language Structures (WALS; Dryer & Haspelmath 2013), Grambank (Skirgård et al., 2023), and the AUTOTYP Database (Bickel et al., 2023). For each of these databases, I conducted a one-way ANOVA with the N1 ratio as the dependent variable, and the database category as the grouping variable for all languages shared between the PBC and the respective database (WALS: 646 languages; Grambank: 937 languages; AUTOTYP: 203 languages). [^30] Code to reproduce these results can be found at the linked repository. For each comparison, the N1 ratio significantly differentiated between SV and VS languages (WALS mean diff. 1.443, p < 0.001; Grambank mean diff. 1.285, p < 0.001; AUTOTYP mean diff. 1.544, p < 0.001), and for WALS and Grambank it additionally differentiated between SV/VS and free word orders, [^31] For AUTOTYP the N1 ratio significantly differentiated between SV and free word orders (p ¡ 0.001) but was marginal for VS and free languages (p = 0.054). indicating that this ratio can be used as a proxy for word order, at least for the PBC.
Figure 2: Distrubution of the N1 ratio in the taggedPBC
@@
This test of features indicates that the taggedPBC is a valid source of crosslinguistic comparative data. Despite potential inaccuracies in tags, features extracted directly from these corpora correlate well with expert tools and annotations, both in terms of tags in individual languages and in terms of features such as word order. This suggests some exciting new lines of research.
@@ 4 Conclusions and future work
The preceding sections have presented a new, massive dataset of parallel text tagged for parts of speech in 1,942 languages via cross-lingual POS-tag transfer, derived partially from the PBC (Mayer & Cysouw 2014) with additional data sourced from newly-accessible Bible translations. The dataset is available for research on Github as the taggedPBC, allowing for expansion/collaboration, and the tags for nouns and verbs have been validated via comparison with taggers for high-resource languages as well as hand-tagged data from the UDT. Additionally, I have shown that the N1 ratio, a feature extracted from this dataset on the basis of arguments and predicates, differentiates between the intransitive word order (SV, VS, free) of languages, allowing for the classification of previously unclassified languages, and suggesting exciting new directions for research in computational typology.
@@ 4.2 Potential research directions
This dataset provides an important source of data for the growing field of computational typology. As mentioned in footnote 9 above, this field has primarily been focused on statistical methods whereby different features of language (or of environment) are correlated with determinations or classifications in typological databases (see Jäger 2025). This kind of investigation is limited by the availability of information on such features or classifications for particular languages, and there may be some debate regarding how to precisely categorize a language (with respect to word order, see the large number of “UNK” [unknown] classifications for word order in the Grambank database).
@@
The dataset presented here, along with the example of extracting the N1 ratio and observing its relation to word order, provides a potential new direction for computational typology, whereby features can be directly extracted from corpora. As with the example above, datasets like the taggedPBC allow researchers to identify corpus-based features that correlate with (and/or differentiate between) expert classifications in typological databases. These features can then be used to classify previously unclassified languages, greatly expanding the coverage of such databases, and directly addressing the issue of linguistic diversity, as highlighted by Ben´ıtez-Burraco (2025) and others. Further, such corpus-derived measures allow for differing approaches to linguistic questions, such as the gradient-based word order approach proposed by Levshina (2023). While the example above shows that languages do not overlap very much in terms of the N1 ratio with regard to basic intransitive word order classifications, the point remains that no language is entirely SV or VS, supporting alternative methods to answer related questions.
@@
Some questions that this dataset is uniquely able to support investigations in are those related to word types crosslinguistically, whether with regard to length, order, or other patterns. The fact that such questions can be queried in relation to corpora additionally allows for known effects of frequency on language (Behrens & Pfänder 2016; Bentz et al. 2017; Bybee 2006; Haspelmath 2021; Koplenig et al. 2022) to be accounted for. Other questions that could be asked, related to dependencies and phonological properties, may require additional annotation, whether automated, semi-automated, or manual. While improvements can definitely be made to the overall annotations, the taggedPBC dataset provides a solid baseline to push computational typology forward as a means of investigating (universal) properties of human language.

### v3

@@ 3 Results
The final taggedPBC dataset contains tagged data for 1,942 modern languages (1,944 if you include the 2 conlangs with NT translations: Esperanto and Klingon). This represents roughly 25% of the world’s languages from 155 distinct families and 78 single languages or isolates. The number of languages with more than 1,800 sentences/verses in the dataset is 1,853 (see Table 1), and the fewest number of verses is represented by Nafusi (ISO 639-3: jbn) with 729 verses. The automatic tagging method outlined above results in a statistically robust number of unique arguments and predicates [^20] In the current paper, I use the term “argument” and “predicate” as a way of grouping classes of words that loosely correspond to heads of phrases and serve similar roles/functions crosslinguistically. Specifically, “argument” groups the word classes of nouns, proper nouns, and pronouns, which are deemed to represent prototypical agents/patients. The term “predicate” groups the word classes of verbs and auxiliaries, which are deemed to represent prototypical actions/states. By using these terms to group POS tags, I am not claiming that these tags exclusively represent all instantiations of “arguments” or “predicates” crosslinguistically, but simply separating out those tags from other tags in the Universal Tag Set used in the overall dataset, in order to develop procedures to compute relevant statistics for crosslinguistic comparison. for all 1,942 languages (mean: 699 arguments, 466 predicates), with the fewest number of unique arguments being tagged for Nugunu (ISO 639-3: yas; 76) and the fewest number of unique predicates being tagged for Rapanui (ISO 639-3: rap; 57). If we compare the number of unique nouns and verbs, the fewest unique nouns are identified for ut-Hun (ISO 639-3: uth; 59) and the fewest unique verbs for Bambalang (ISO 639-3: bmo; 54).
@@ 3.1 Validating tag accuracy
One way to validate the automated tags provided by the IBM2 word alignment method involves assessing against other automatic taggers developed for high resource languages. Here we want to see that the method used for tagging these languages aligns with existing language-specific taggers at a level above chance.
POS-taggers have been developed for high-resource languages, and the current state of the art is represented by the SpaCy and Trankit text processing libraries. These libraries have supervised POS-taggers for 16 and 31 languages respectively (excluding languages with non-roman scripts). For this comparison, each of the trained taggers in these libraries was used to tag the reduced PBC verses for the respective language and then the specific tags for arguments (NOUN, PROPN, PRON) and predicates (VERB, AUX) were compared with the same verses tagged via the word alignment method. [^21] I should note here, first, that both libraries use a variation of the universal tag set adopted by the UDT project and my methodology, as outlined in footnote 19. While there are some differences, they apply to less common word classes and do not affect arguments and predicates, which were the focus of this comparison. Second, these taggers were not developed using Biblical texts, which reduces the overall likelihood of correspondence, and makes the actual result more striking. Specifically, each word-alignment tagged verse was stripped of tags and re-tagged by the respective libraries. Then the two sets of tagged verses (between the library and the word-alignment method) were compared and the number of matching arguments/predicates was divided by the total number of arguments/predicates identified by the respective library in a given sentence. This was then averaged across all verses to provide an agreement percentage between the translation method and the respective library for the argument and predicate tags. [^22] See code at the linked repository for a more detailed breakdown of correspondence between individual POS tags in the argument and predicate classes. Correspondence with argument and predicate tags between the word alignment method and these pos-taggers was on average above 75% for both SpaCy (Table 2; arguments: 0.84; predicates: 0.76) and Trankit (Table 3; arguments: 0.84; predicates: 0.77).
@@
This comparison informs us that the translation method agrees with existing SOTA POS-taggers more than 75% of the time, which suggests that this methodology is relatively robust at transferring tags from English to other languages. At least for tagging major word classes, it aligns pretty well with automatic taggers individually developed for specific languages. However, “the proof is in the pudding”: does the data in the taggedPBC lend itself to actual correlations with observed properties of language? To assess this, we now turn to observations of word order that can be extracted from this dataset.
@@ 3.2 Testing features extracted from the dataset: word order
By “word order” most researchers refer to the relative order of referential arguments (i.e. elements referring to prototypical agents/patients, often represented by nouns, proper nouns, and pronouns) versus predicates (i.e. elements referring to prototypical actions and states as represented by verbs and auxiliaries). This makes it possible to talk about “verb-initial” vs “verb-final” or “verb-medial” languages, as well as grammatical constructs like Subject and Object in relation to the position of the Verb. In the current investigation I focus on intransitive word order, which is concerned with a predicate that requires a single core argument, resulting in three possible word orders: “verb final” (SV), “verb initial” (VS), and “free” (no dominant order). In order to identify these relations in corpora I use the following procedure for each language in the taggedPBC corpus:
@@
    Filter the verses in the corpus, retaining only verses/sentences with both an argument and a predicate.
@@
    In the resulting set, verses/sentences with an argument occurring before a predicate are classed as “noun first”, while those with a predicate occurring before an argument are classed as “verb first”. [^23] Note that the terms “noun first” and “verb first” are simply class/grouping labels (they could be termed “X” and “Y”) used to compute the N1 ratio. The terms do not constitute claims about a particular sentence structure.
@@
    This gives a count of “noun first” and “verb first” verses/sentences for each language.
@@
    Turn this into the “N1 ratio” by dividing the total number of “noun first” verses/sentences by the total number of “verb first” verses/sentences.
@@
To test the validity of this measure I compared the derived values with known expert judgements of intransitive word order (SV, VS, or “free”) in three typological databases (Figure 1): the World Atlas of Language Structures (WALS; [Dryer:2013ab]), Grambank [Skirgard:2023ab], and the AUTOTYP Database [Bickel:2023aa]. For each of these databases, I conducted a one-way ANOVA with the N1 ratio as the dependent variable, and the database category as the grouping variable for all languages shared between the PBC and the respective database (WALS: 646 languages; Grambank: 937 languages; AUTOTYP: 203 languages). [^24] Code to reproduce these results can be found at the linked repository. For each comparison, the N1 ratio significantly differentiated between SV and VS languages (WALS mean diff. 1.443, p < 0.001; Grambank mean diff. 1.285, p < 0.001; AUTOTYP mean diff. 1.544, p < 0.001), and for WALS and Grambank it additionally differentiated between SV/VS and free word orders, [^25] For AUTOTYP the N1 ratio significantly differentiated between SV and free word orders (p < 0.001) but was marginal for VS and free languages (p = 0.054). indicating that this ratio can be used as a proxy for word order, at least for the PBC.
Figure 2: Distribution of the N1 ratio in the taggedPBC
@@
This test of features indicates that the taggedPBC is a valid source of crosslinguistic comparative data. Despite potential inaccuracies in tags, features extracted directly from these corpora correlate well with expert tools and classifications, both in terms of tags in individual languages and in terms of features such as word order. This suggests some exciting new lines of research.
@@ 4 Conclusions and future work
The preceding sections have presented a new, massive dataset of parallel text tagged for parts of speech in 1,942 languages via cross-lingual POS-tag transfer, derived partially from the PBC [Mayer:2014aa] with additional data sourced from newly-accessible Bible translations. The dataset is available for research on Github as the taggedPBC, allowing for expansion/collaboration, and the tags for nouns and verbs have been validated via comparison with taggers for high-resource languages. Additionally, I have shown that the N1 ratio, a feature extracted from this dataset on the basis of arguments and predicates, differentiates between the intransitive word order (SV, VS, free) of languages, allowing for the classification of previously unclassified languages, and suggesting exciting new directions for research in computational typology.
@@ 4.2 Potential research directions
This dataset provides an important source of data for the growing field of computational typology. As mentioned in footnote 4 above, this field has primarily been focused on statistical methods whereby different features of language (or of environment) are correlated with determinations or classifications in typological databases (see [Jager:2025aa]). This kind of investigation is limited by the availability of information on such features or classifications for particular languages, and there may be some debate regarding how to precisely categorize a language (with respect to word order, see the large number of “UNK” [unknown] classifications for word order in the Grambank database).
@@
The dataset presented here, along with the example of extracting the N1 ratio and observing its relation to word order, provides a potential new direction for computational typology, whereby features can be directly extracted from corpora. As with the example above, datasets like the taggedPBC allow researchers to identify corpus-based features that correlate with (and/or differentiate between) expert classifications in typological databases. These features can then be used to classify previously unclassified languages, greatly expanding the coverage of such databases, and directly addressing the issue of linguistic diversity, as highlighted by [Benitez-Burraco:2025aa] and others. Further, corpus-derived measures allow for differing approaches to linguistic questions, such as the gradient-based word order approach proposed by [Levshina:2023ab]. While the example above shows that languages do not overlap very much in terms of the N1 ratio with regard to basic intransitive word order classifications, the point remains that no language is entirely SV or VS, supporting alternative methods to answer related questions.
In general, if we consider POS tags to be one means of identifying syntactic patterns, as suggested by the present investigation, there should be a principled, crosslinguistic way to do so. Grouping POS tags and/or observing their relative orders in sentences is one way to compute measures from corpora. It logically follows, then, that if such measures reliably correlate with other linguistic features (as the N1 ratio does with intransitive word order in WALS/Grambank/Autotyp), we should be able to use them as proxy measures for those features. Having a tagged crosslinguistic dataset of this size is valuable for answering multiple questions (such as whether a language has prepositions or postpositions), and improving the scope and quality of the annotations simply increases the kinds of questions that can be asked and answered.
@@
Some questions that this dataset is uniquely able to support investigations in are those related to word types crosslinguistically, whether with regard to length, order, or other patterns. The fact that such questions can be queried in relation to corpora additionally allows for known effects of frequency on language [Behrens:2016aa, Bentz:2017aa, Bybee:2006aa, Haspelmath:2021aa, Koplenig:2022aa] to be accounted for. Other questions that could be asked, related to dependencies and phonological properties, may require additional annotation, whether automated, semi-automated, or manual. While improvements can definitely be made to the overall annotations, the taggedPBC dataset provides a solid baseline to push computational typology forward as a means of investigating (universal) properties of human language.